I think the "= 0" is messing me up...
(x+6)(2x-1) = 0
I tried:
(1x+6)(2x-1) = 0
2x - 1x + 12x - 6 = 0
2x (squared) - 11x - 6 = 0
and now I'm stuck, I'm supposed to get 2 answers? Thanks.

Answer by Earlsdon(6294)

You might want to review the "zero products principle" which states:
If a*b = 0, then either a = 0, or b = 0, or both.
Apply this principle to your problem:
If (x+6)(2x-1) = 0, then either (x+6) = 0, or (2x-1) = 0, or both.
If (x+6) = 0, then x = -6
If (2x-1) = 0, then x = 1/2
The two solutions are:
x = -6
x = 1/2